OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

OpenStudy (cwrw238):

cos x = u . v / |u| |v| where u.v is the dot product and |u| and |v| are their magnitudes

OpenStudy (anonymous):

So cosx = (8)(9) + (4)(-9) / (1)(15) = 72 + 36 = 108 / 15 = 7.2 Right @cwrw238 ?

OpenStudy (cwrw238):

dot product is 72 - 36 = 36

OpenStudy (cwrw238):

magnitudes are sqrt(8^2 + 4^2) and sqrt (9^2 + (-9)^2)

OpenStudy (anonymous):

36 / sqrt(64+16) * sqrt(81 + 81) 36/ 4sqrt5*9sqrt2

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

idk what to do next man :l

OpenStudy (cwrw238):

i get that = 0.3162 cos x = 0.3162

OpenStudy (cwrw238):

just enter this value into your calculator and press cos-1 key to get x x is the required angle

OpenStudy (anonymous):

Got my answer! thanks cwrw

OpenStudy (cwrw238):

yw